Hi
If Simon gave 30% of his sweets to Terry they will have the same number of sweets. If Simon gave 250 sweets to Terry, Terry will 80% more sweets than Simon. How many sweets does Simon have.

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52781)

.
If Simon gave 30% of his sweets to Terry they will have the same number of sweets.
If Simon gave 250 sweets to Terry, Terry will 80% more sweets than Simon.
How many sweets does Simon have.
~~~~~~~~~~~~~~~~~~~~~~

x = # of Simon' sweets;

y = # of Terry' sweets.


From the problem, we have two equations.

First equation is

    (1-0.3)x = y + 0.3x   (first statement),

which  simplifies to

    0.7x = y + 0.3x  --->  0.7x - 0.3x = y  --->  0.4x = y.



Second equation is

    x - 250 = = (y + 250) - 0.8*(x-250)  (second statement),

which  simplifies to

    x - 250 = y + 250  - 0.8x + 200,

    x - 250 - 250 - 200 = y - 0.8x

    x - 700 = y - 0.8x

    y = 1.8x - 700.

    

So, we have this system of two equations

    y = 0.4x             (1)

    y = 1.8x - 700       (2)


Equations  (1) and (2) have left sides identical, so their right sides are equal

    0.4x = 1.8x - 700

    700 = 1.4x 

    x = 700/1.4 = 500.


Thus Simon has 500 sweets.

From equation (1), Terry has 0.4*x = 0.4*500 = 200 sweets.


ANSWER.  Simon has 500 sweets.

Solved.



Answer by math_tutor2020(3817)


Answer: 500

Explanation

x = number of sweets Simon starts with
y = number of sweets Terry starts with

Scenario 1. Simon gives 30% of his sweets to Terry. After this they end up with the same number of sweets.

Scenario 2. Simon gives 250 sweets to Terry. Afterward, Terry will have 80% more sweets compared to Simon.

Scenario 1BeforeAfter
Simonx0.7x
Terryyy+0.3x

Since they end up with the same number of sweets, we can say 0.7x = y+0.3x
Solve for y to get y = 0.4x
We'll use this later.

Scenario 2BeforeAfter
Simonxx-250
Terryyy+250

At the end of this scenario, Terry has 80% more sweets compared to Simon.
Think of it like this
TerrysCount = SimonsCount + 80% of SimonsCount
TerrysCount = SimonsCount + 0.8*SimonsCount
TerrysCount = (1 + 0.8)*SimonsCount
TerrysCount = 1.8*SimonsCount
Or think of "Terry has 80% more" as "Terry's count is 180% of Simon's count". 180% then converts to the decimal form 1.8

Based on that template, we can then say,
TerrysCount = 1.8*SimonsCount
y+250 = 1.8*(x-250)
0.4x+250 = 1.8*(x-250) .............. plug in y = 0.4x
0.4x+250 = 1.8x-450
1.8x-0.4x = 250+450
1.4x = 700
x = 700/1.4
x = 500
Simon started with 500 sweets.
y = 0.4x = 0.4*500 = 200 is the number of sweets Terry started with.

--------------------------------------

Let's check the answer.
To do that we go through each scenario.
I'll update each previous table to plug in the mentioned x and y values.
Scenario 1BeforeAfter
Simon500350
Terry200350

Both end up with an equal number of candies (each with 350).
This confirms scenario 1.

Scenario 2BeforeAfter
Simon500250
Terry200450

Divide Terry's new count over Simon's new count to get 450/250 = 1.8 to show that Terry has 80% more compared to Simon.
Or you can notice that Terry has 450-250 = 200 more candies compared to Terry, and 200/250 = 0.80 = 80%

Both scenarios are confirmed. This confirms the answer.